# xLSTMAD: A Powerful xLSTM-based Method for Anomaly Detection

### v1

@@ IV-B Reproducibility guidelines
  • xLSTMAD-R: window size: {15, 25, 50}, learning rate: {$0.005$, $0.001$, $0.0008$}, embedding dimensionality: {20, 40}.
  • xLSTMAD-F: window size: {15, 25, 50}, learning rate: {$0.005$, $0.001$, $0.0008$}, embedding dimensionality: {20, 40}, forecasting horizon: {1, 5, 15}.
@@ V Results Discussion
LSTMAD [3] 0.31 0.74 0.31 0.7 0.36 0.79 0.64 0.38 0.87
USAD [20] 0.3 0.68 0.26 0.64 0.31 0.53 0.4 0.37 0.8
AutoEncoder [12] 0.3 0.69 0.3 0.67 0.34 0.6 0.44 0.28 0.8
KMeansAD [21] 0.29 0.73 0.25 0.69 0.31 0.68 0.49 0.33 0.82
CBLOF [22] 0.27 0.7 0.28 0.67 0.32 0.65 0.45 0.31 0.81
MCD [23] 0.27 0.69 0.27 0.65 0.33 0.46 0.33 0.2 0.76
OCSVM [24] 0.26 0.67 0.23 0.61 0.28 0.48 0.41 0.3 0.8
Donut [25] 0.26 0.71 0.2 0.64 0.28 0.52 0.36 0.21 0.81
RobustPCA [26] 0.24 0.61 0.24 0.58 0.29 0.6 0.42 0.33 0.81
FITS [27] 0.21 0.66 0.15 0.58 0.22 0.72 0.32 0.16 0.81
OFA [28] 0.21 0.63 0.15 0.55 0.21 0.72 0.41 0.17 0.83
EIF [29] 0.21 0.71 0.19 0.67 0.26 0.74 0.44 0.26 0.81
COPOD [11] 0.2 0.69 0.2 0.65 0.27 0.72 0.41 0.24 0.8
IForest [30] 0.2 0.69 0.19 0.66 0.26 0.68 0.41 0.24 0.8
HBOS [31] 0.19 0.67 0.16 0.63 0.24 0.67 0.4 0.24 0.8
TimesNet [32] 0.19 0.64 0.13 0.56 0.2 0.68 0.32 0.17 0.82
KNN [33] 0.18 0.59 0.14 0.51 0.19 0.69 0.45 0.21 0.79
TranAD [4] 0.18 0.65 0.14 0.59 0.21 0.68 0.4 0.21 0.79
LOF [34] 0.14 0.6 0.1 0.53 0.15 0.57 0.32 0.14 0.76
AnomalyTransformer [5] 0.12 0.57 0.07 0.52 0.12 0.53 0.33 0.14 0.74
@@
CATSv2 [35] 0.10 0.10 0.04 0.04 0.04 0.08 0.03 0.61 0.61 0.46 0.46 0.43 0.64 0.53
CreditCard [36] 0.06 0.07 0.02 0.02 0.02 0.02 0.02 0.75 0.81 0.87 0.87 0.87 0.86 0.84
Daphnet [37] 0.50 0.46 0.31 0.31 0.31 0.21 0.06 0.95 0.95 0.81 0.81 0.81 0.84 0.52
Exathlon [38] 0.91 0.91 0.82 0.82 0.82 0.68 0.10 0.95 0.95 0.89 0.89 0.89 0.88 0.51
GECCO [39] 0.04 0.04 0.02 0.02 0.02 0.03 0.02 0.50 0.54 0.44 0.44 0.44 0.65 0.64
GHL [40] 0.01 0.01 0.16 0.16 0.06 0.02 0.01 0.46 0.45 0.69 0.68 0.63 0.55 0.52
Genesis [41] 0.01 0.01 0.03 0.03 0.04 0.10 0.01 0.78 0.85 0.58 0.58 0.58 0.96 0.77
LTDB [42] 0.40 0.34 0.34 0.34 0.31 0.33 0.19 0.70 0.68 0.71 0.71 0.71 0.72 0.58
MITDB [42] 0.09 0.08 0.13 0.13 0.11 0.14 0.04 0.66 0.66 0.69 0.68 0.67 0.69 0.54
MSL [43] 0.38 0.40 0.41 0.34 0.23 0.35 0.08 0.81 0.81 0.81 0.77 0.73 0.77 0.63
OPPORT. [44] 0.16 0.16 0.16 0.16 0.17 0.16 0.05 0.28 0.28 0.65 0.65 0.65 0.61 0.53
PSM [45] 0.18 0.18 0.24 0.24 0.22 0.22 0.13 0.61 0.62 0.74 0.73 0.72 0.70 0.54
SMAP [43] 0.35 0.32 0.20 0.22 0.17 0.20 0.04 0.73 0.74 0.76 0.76 0.71 0.78 0.60
SMD [13] 0.35 0.37 0.31 0.30 0.31 0.37 0.05 0.83 0.84 0.83 0.83 0.83 0.83 0.60
SVDB [46] 0.26 0.19 0.20 0.20 0.15 0.19 0.06 0.68 0.67 0.72 0.72 0.69 0.73 0.57
SWaT [47] 0.34 0.35 0.16 0.16 0.16 0.48 0.14 0.71 0.71 0.50 0.50 0.47 0.74 0.53
TAO [48] 0.80 0.82 1.00 1.00 0.99 1.00 0.77 0.88 0.91 1.00 1.00 1.00 1.00 0.94
@@
In the following discussions, we focus on comparisons between our proposed method and two baselines: CNN [2] and LSTMAD [3]. While CNN is relevant since it achieved the best performance in the TSB-AD benchmark [1], LSTMAD is a natural competitor since it is based on LSTM, the predecessor of xLSTM. Moreover, to support the interpretation of experimental results, we provide a Random model that simply returns a random anomaly score in the range from 0 to 1. This allows us to set realistic expectations for VUS-based metric values for non-random anomaly detection methods.
@@ V-B Fine-grained anomaly detection results
On Daphnet, xLSTMAD-R (MSE) achieves a VUS-PR of 0.50 compared to 0.31 for LSTM and 0.21 for CNN. On high-noise or sparse anomaly datasets such as CATSv2 and GECCO, the performance margins are smaller, though xLSTMAD-R (SoftDTW) still outperforms CNN and LSTM in most cases. On the other hand, all variants of xLSTMAD achieve very low VUS-PR results for the Genesis dataset. The reason behind such results may be the fact that the Genesis dataset consists of just three very short anomalies [1]. The TAO dataset is a notable example with perfect or near-perfect scores across all models. All four variants of our approach achieve VUS-ROC = 1.00 and two of them achieve VUS-PR = 1.00 xLSTMAD-F, indicating the ease of detecting anomalies in this scenario.

### v2

@@ IV-C Reproducibility guidelines
  • xLSTMAD-R: window size: {15, 25, 50}, learning rate: {$0.005$, $0.001$, 0.00080.0008}, embedding dimensionality: {20, 40}.
  • xLSTMAD-F: window size: {15, 25, 50}, learning rate: {$0.005$, $0.001$, 0.00080.0008}, embedding dimensionality: {20, 40}, forecasting horizon: {1, 5, 15}.
@@ V Results Discussion
  • RQ3. Are there sensible differences in predictive behavior between the proposed anomaly detection losses (MSE vs. SoftDTW) and anomaly detection approaches (xLSTMAD-F vs. xLSTMAD-R)?
@@
LSTMAD [ji2021novel] 0.31 0.74 0.31 0.7 0.36 0.79 0.64 0.38 0.87
USAD [audibert2020usad] 0.3 0.68 0.26 0.64 0.31 0.53 0.4 0.37 0.8
AutoEncoder [sakurada2014anomaly] 0.3 0.69 0.3 0.67 0.34 0.6 0.44 0.28 0.8
KMeansAD [yairi2001fault] 0.29 0.73 0.25 0.69 0.31 0.68 0.49 0.33 0.82
CBLOF [he2003discovering] 0.27 0.7 0.28 0.67 0.32 0.65 0.45 0.31 0.81
MCD [estimator1999fast] 0.27 0.69 0.27 0.65 0.33 0.46 0.33 0.2 0.76
OCSVM [hejazi2013one] 0.26 0.67 0.23 0.61 0.28 0.48 0.41 0.3 0.8
Donut [xu2018unsupervised] 0.26 0.71 0.2 0.64 0.28 0.52 0.36 0.21 0.81
RobustPCA [paffenroth2018robust] 0.24 0.61 0.24 0.58 0.29 0.6 0.42 0.33 0.81
FITS [xufits] 0.21 0.66 0.15 0.58 0.22 0.72 0.32 0.16 0.81
OFA [zhou2023one] 0.21 0.63 0.15 0.55 0.21 0.72 0.41 0.17 0.83
EIF [hariri2019extended] 0.21 0.71 0.19 0.67 0.26 0.74 0.44 0.26 0.81
COPOD [li2020copod] 0.2 0.69 0.2 0.65 0.27 0.72 0.41 0.24 0.8
IForest [liu2008isolation] 0.2 0.69 0.19 0.66 0.26 0.68 0.41 0.24 0.8
HBOS [goldstein2012histogram] 0.19 0.67 0.16 0.63 0.24 0.67 0.4 0.24 0.8
TimesNet [wutimesnet] 0.19 0.64 0.13 0.56 0.2 0.68 0.32 0.17 0.82
KNN [ramaswamy2000efficient] 0.18 0.59 0.14 0.51 0.19 0.69 0.45 0.21 0.79
TranAD [tuli2022tranad] 0.18 0.65 0.14 0.59 0.21 0.68 0.4 0.21 0.79
LOF [breunig2000lof] 0.14 0.6 0.1 0.53 0.15 0.57 0.32 0.14 0.76
AnomalyTransformer [xuanomaly] 0.12 0.57 0.07 0.52 0.12 0.53 0.33 0.14 0.74
@@
CATSv2 [CATSV2] 0.10 0.10 0.04 0.04 0.04 0.08 0.03 0.61 0.61 0.46 0.46 0.43 0.64 0.53
CreditCard [sharafaldin2018toward] 0.06 0.07 0.02 0.02 0.02 0.02 0.02 0.75 0.81 0.87 0.87 0.87 0.86 0.84
Daphnet [bachlin2009wearable] 0.50 0.46 0.31 0.31 0.31 0.21 0.06 0.95 0.95 0.81 0.81 0.81 0.84 0.52
Exathlon [jacob2020exathlon] 0.91 0.91 0.82 0.82 0.82 0.68 0.10 0.95 0.95 0.89 0.89 0.89 0.88 0.51
GECCO [moritz2018gecco] 0.04 0.04 0.02 0.02 0.02 0.03 0.02 0.50 0.54 0.44 0.44 0.44 0.65 0.64
GHL [filonov2016multivariate] 0.01 0.01 0.16 0.16 0.06 0.02 0.01 0.46 0.45 0.69 0.68 0.63 0.55 0.52
Genesis [von2018anomaly] 0.01 0.01 0.03 0.03 0.04 0.10 0.01 0.78 0.85 0.58 0.58 0.58 0.96 0.77
LTDB [goldberger2000physiobank] 0.40 0.34 0.34 0.34 0.31 0.33 0.19 0.70 0.68 0.71 0.71 0.71 0.72 0.58
MITDB [goldberger2000physiobank] 0.09 0.08 0.13 0.13 0.11 0.14 0.04 0.66 0.66 0.69 0.68 0.67 0.69 0.54
MSL [hundman2018detecting] 0.38 0.40 0.41 0.34 0.23 0.35 0.08 0.81 0.81 0.81 0.77 0.73 0.77 0.63
OPPORT. [roggen2010collecting] 0.16 0.16 0.16 0.16 0.17 0.16 0.05 0.28 0.28 0.65 0.65 0.65 0.61 0.53
PSM [abdulaal2021practical] 0.18 0.18 0.24 0.24 0.22 0.22 0.13 0.61 0.62 0.74 0.73 0.72 0.70 0.54
SMAP [hundman2018detecting] 0.35 0.32 0.20 0.22 0.17 0.20 0.04 0.73 0.74 0.76 0.76 0.71 0.78 0.60
SMD [su2019robust] 0.35 0.37 0.31 0.30 0.31 0.37 0.05 0.83 0.84 0.83 0.83 0.83 0.83 0.60
SVDB [greenwald1990improved] 0.26 0.19 0.20 0.20 0.15 0.19 0.06 0.68 0.67 0.72 0.72 0.69 0.73 0.57
SWaT [mathur2016swat] 0.34 0.35 0.16 0.16 0.16 0.48 0.14 0.71 0.71 0.50 0.50 0.47 0.74 0.53
TAO [TAO] 0.80 0.82 1.00 1.00 0.99 1.00 0.77 0.88 0.91 1.00 1.00 1.00 1.00 0.94
@@
In the following discussions, we focus on comparisons between our proposed method and two baselines: CNN [munir2018deepant] and LSTMAD [ji2021novel]. While CNN is relevant since it achieved the best performance in the TSB-AD benchmark [liu2024elephant], LSTMAD is a natural competitor since it is based on LSTM, the predecessor of xLSTM. Moreover, to support the interpretation of experimental results, we provide a Random model that simply returns a random anomaly score in the range from 0 to 1. This allows us to set realistic expectations for VUS-based metric values for non-random anomaly detection methods.
@@ V-B Fine-grained anomaly detection results
On Daphnet, xLSTMAD-R (MSE) achieves a VUS-PR of 0.50 compared to 0.31 for LSTM and 0.21 for CNN. On high-noise or sparse anomaly datasets such as CATSv2 and GECCO, the performance margins are smaller, though xLSTMAD-R (SoftDTW) still outperforms CNN and LSTM in most cases. On the other hand, all variants of xLSTMAD achieve very low VUS-PR results for the Genesis dataset. The reason behind such results may be the fact that the Genesis dataset consists of just three very short anomalies [liu2024elephant]. The TAO dataset is a notable example with perfect or near-perfect scores across all models. All four variants of our approach achieve VUS-ROC = 1.00 and two of them achieve VUS-PR = 1.00 xLSTMAD-F, indicating the ease of detecting anomalies in this scenario.
@@
V-C Comparison of xLSTMAD variants
A closer inspection of the four xLSTMAD variants reveals distinct behavioral patterns.
Forecasting vs. Reconstruction modes: Based on results in Table II, the reconstruction-based variant xLSTMAD-R leads in average performance according to VUS-PR. However, the forecasting-based variant xLSTMAD-F outperforms xLSTMAD-R across nearly all other metrics, showing complementary strengths, especially when temporal alignment is critical. For example, xLSTMAD-F (MSE) achieves 0.70 Event-based-F1 vs. 0.45 for xLSTMAD-R (MSE), and 0.89 Affiliation-F1 vs. 0.82, suggesting that forecasting-based anomaly detection is more effective at modeling latent temporal expectations than reconstruction-based approaches.
Loss Type (MSE vs. SoftDTW): The SoftDTW-based loss offers competitive gains, particularly in reconstruction mode in terms of VUS-ROC, for some datasets such as CreditCard (from 0.75 to 0.81), GECCO (from 0.50 to 0.54), and TAO (from 0.88 to 0.91), due to their ability to accommodate small temporal shifts and warping in sequences. However, it can be observed that the use of SoftDTW does not consistently outperform MSE. From a more global perspective, for both forecasting and reconstruction modes, the MSE variant exhibits marginally better results across most metrics, such as AUC-PR and Event-based-F1. This indicates that while SoftDTW introduces alignment-aware sensitivity, it may also increase noise sensitivity or reduce stability during training.
In summary, these comparisons emphasize that both model design and loss function selection play crucial roles in multivariate anomaly detection performance, with xLSTMAD-F and MSE emerging as the most robust combination on the TSB-AD-M benchmark (RQ3).
V-D Empirical time complexity
We analyze the average execution time (in seconds, using Nvidia GH200 GPU) of different anomaly detection methods on all 180 time series originating from 17 datasets (including training and testing steps). As expected, reconstruction-based approaches incur the highest computational cost of 626 and 355 seconds on average for MSE and SoftDTW losses, respectively. This computational cost stems mostly from the complexity of reconstructing multivariate sequences (window size of 50 and 25 time steps, respectively, see Table I). Prediction-based models offer better efficiency (121 seconds for MSE loss and 264 seconds for SoftDTW loss). This improvement is related to the fact that instead of reconstructing the whole window, the model predicts only time steps from a much shorter forecasting horizon (5 time steps in our experiments).
In contrast, baseline deep learning models like LSTM and CNN are significantly faster, with CNN achieving the lowest execution time of just 17.12 seconds. This highlights a clear trade-off between detection performance and computational efficiency: while xLSTM-based methods (especially those using DTW) provide robust anomaly detection performance, they do so at a higher computational cost. However, it should be noted that xLSTM models are still in their infancy, and improvements in their efficiency may be achieved with future implementations that simplify the architecture or leverage hardware acceleration more extensively.
@@ VI Conclusion
Acknowledgments
Project cofunded by: i) the Polish Ministry of Science and Higher Education under the program ”Co-funded International Projects,”; ii) the program ”Excellence initiative-research university” for the AGH University in Krakow; iii) the ARTIQ project: UMO-2021/01/2/ST6/00004 and ARTIQ/0004/2021.